What is the magnitude and direction (relative to the +a axis) of the net electric force due to the other three charges on qi in the upper-left corner of the rectangle below? +3 42 q1 91 = +1 C 5 m 3 m 43 = 36.90 4 m
Answer #1

The force on charge q1 due to q3 which will be given as -

F13 = ke |q1| |q3| / r132

F13 = [(9 x 109 Nm2/C2) (1 C) (3 C)] / (5 m)2

F13 = 1.08 x 109 N

The x and y components of F13 which are given below as -

F13,x = (1.08 x 109 N) cos 36.90

F13,x = 8.63 x 108 N

And

F13,y = (1.08 x 109 N) sin 36.90

F13,y = 6.48 x 108 N

The force on charge q1 due to q2 which will be given as -

F12 = ke |q1| |q2| / r122

F12 = [(9 x 109 Nm2/C2) (1 C) (2 C)] / (4 m)2

F12 = 1.125 x 109 N

The x and y components of F12 which are given below as -

F12,x = (1.125 x 109 N) cos 00

F12,x = 1.125 x 109 N

And

F12,y = (1.125 x 109 N) sin 00

F12,y = 0 N

The net electric force due to other three charges on q1 which will be given as -

F1 = (F13,x + F12,x) + (F13,y + F12,y)

F1 = [(8.63 x 108 N) + (1.125 x 109 N)] + [(6.48 x 108 N) + (0 N)]

F1 = (1.98 x 109 N) + (6.48 x 108 N)

Therefore, magnitude of the net electric force due to other three charges on q1 which will be given by -

| F1 | = \sqrt{}(1.98 x 109 N)2 + (6.48 x 108 N)2

| F1 | = \sqrt{}4.34 x 1018 N2

| F1 | = 2.08 x 109 N

From a trigonometry identity, we have

\phi = tan-1 [(6.48 x 108 N) / (1.98 x 109 N)]

\phi = tan-1 (0.3272)

\phi = 18.1 degree